OpenStudy (anonymous):

A wholesaler sells 20 pens for the marked price of 16 pens to a retailer. The retailer in turn sells them at marked price. The retailer a) makes a profit of 25%. b) suffers a loss of 25%. c) makes a profit of 15%. d) none of the above.

OpenStudy (anonymous):

a) (20-16)/16 = 1/4 = 25%

OpenStudy (anonymous):

Thanks, but could you explain a little more elaborately?

OpenStudy (anonymous):

say it's a dollar per pen, he makes $20, and paid $16, so the profit is 20-16=$4 as a percentage of his original cost, 4/16*100=25%

OpenStudy (anonymous):

ok, thanks